if 3/8 of an inch is proportional to 15 inches, how many inches are proportional to six feet
Answer by tommyt3rd(5050)

3/8 : 15 = x : 72
(3/8)/15=x/72
1/40=x/72
x=72/40
x=9/5
x=1 4/5 in